A ball is thrown along the vertical. The coordinate of the ball is observed as a function of time y(t). We choose y = 0 at the ground. The ball is at y = 13.2 m at time t = 0, and at y = 5.0 m at time t = 3.2 s. a) Find the displacement and average velocity of the ball between t = 0 and t = 3.2 s. b) Find the initial and final velocity of the ball. c) What is the height of the ball’s peak above ground? d) What is the average speed of the ball between t = 0 and t = 3.2 s?



1
Expert's answer
2021-02-18T18:44:41-0500

The ball is thrown with a velocity "u" from "y=13.2\ m" at "t=0" and it reaches to "y=5\ m" at "t=3.2\ m" .

"a)"

Displacement of ball "(s)= (5-13.2)\ m=-8.2\ m"

Negative sign indicates the displacement in downward direction.

Average velocity of ball "(v_{av})=\frac{-8.2}{3.2}=-2.5625\ m"


"b)"

"a=-9.8\ m/sec^2" (In downward direction )

Using equation "s=ut+\frac{1}{2}at^2" and substituting the values,

"-8.2=u\times 3.2 +\frac{1}{2}(-9.8)(3.2)^2"

"\implies u=+13.1175\ m/sec" which means ball is thrown in vertically upward direction.

Using "v=u+at" and substituting the values,

"v=13.1175+(-9.8)(3.2)"

"\implies v=-18.2425\ m/sec" which means ball is moving downwards.


"c)"

At maximum height , "v=0"

Using equation "v^2-u^2=2as" and substituting the values,

"0^2-13.1175^2=2(-9.8)s"

"\implies s=8.79\ m"


"d)" Average speed "= |v_{av}|=|-2.5625|\ m/sec= 2.5625\ m/sec"